Write expression of Ly in linear momentum form.
1
Expert's answer
2020-11-17T13:04:21-0500
"L_y=zP_x-xP_z=−i\hbar z\frac{\partial }{\partial x}+i\hbar x\frac{\partial }{\partial z}"